Find the value of

 Sin78 - Sin66 -Sin42 + Sin60


all angles are in degree.

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:Hello student, please find answer to your question
V = sin(78)-sin(66)-sin(42)+sin(60)
V = sin(60+18)-sin(60-18)-sin(66)+sin(60)
sin(a+b) = sin(a).cos(b)+cos(a).sin(b)
sin(a-b) = sin(a).cos(b)-cos(a).sin(b)
V = sin(60).cos(18)+cos(60).sin(18)-(sin(60).cos(18)-cos(60).sin(18))-sin(66)+sin(60)V = 2cos(60).sin(18)-sin(66)+sin(60)
V = 2.\frac{1}{2}.sin(18)-sin(66)+sin(60)
V = sin(18)-sin(66)+sin(60)
V = \frac{\sqrt{5}-1}{4}-cos(24)+\frac{\sqrt{3}}{2}